Feeling queasy? Then compute the first three iterates of Z2+i for the following starting complex number z: 5+6i and guess whether the iterates of this number are periodic or tend to infinity if you continued the iteration process. please type

First iterate:

second iterate

third iterate  
Answer #1

Hello,

I have tried to solve the question and write the solution as detailed as I can.

The numerical answers are-

First iterate: (-11)+6i

Second iterate: 3600-(1341)i

Third iterate: 1,11,61,719-(96,55,199)i

When the iterations are repeated, the iterates are not periodic. The sequence tends to infinity. Please check the attached images for the detailed solutions.

Given function is f (2) = z2 ti.. Starting complex number z = 5+ 61 ) f(x) = f(5+61) = (5+61)2 +i . f2) = 5 + 2x5X61 + (61)2Pace No Date is negative and much lesser than the previous iterations This implies that if we continue the iterations, then R